The intensity of sunlight at at the earth is 1.0 kW/m^2. What is the intensity of sunlight at •Mars, which is 1.5 times as far away from the sun?
Answer #1

Here,

for the intensity of the light sources

Intensity = Power/(4pi * r^2)

hence , the intensity is proporational to 1/distance^2

I1/I2 = d2^2/d1^2

I1/1 = 1/1.5^2

I1 = 0.444 kW/m^2 = 444 W/m^2

the intensity of the sunlight at Mars is 444 W/m^2